# Differentially Private Synthetic Data via APIs 3: Using Simulators Instead of Foundation Models

### v1

@@ Appendix B Experimental Details
B.1 MNIST with Text Rendering Program
Tables 5 and 6 show the list of the parameters and their associated feasible sets and variation degrees in the MNIST with Text Rendering Program experiments. The total number of PE iterations is 4.
@@ B.1 MNIST with Text Rendering Program
B.2 CelebA with Generated Images from Computer Graphics-based Render
The variation degrees $\gamma$ across PE iterations are [1000, 500, 200, 100, 50, 20]. The total number of PE iterations is 6.
B.3 CelebA with Rule-based Avatar Generator
@@
These are taken from the input parameters to the library (Escartín, 2021). There is no numerical parameter.
For the experiments with only the simulator, the variation degrees $\beta$ across PE iterations are [0.8, 0.6, 0.4, 0.2, 0.1, 0.08, 0.06]. The total number of PE iterations is 7.
For the experiments with both foundation models and the simulator, we use a total of 5 PE iterations so as to be consistent with the setting in Gong et al. (2025). For the RANDOM_API and the first PE iteration, we use the simulator ($\beta=0.8$). For the next 4 PE iterations, we use the same foundation model as in Lin et al. (2023) with variation degrees [96, 94, 92, 90].

### v2

@@ Appendix I Experimental Details
I.1 MNIST with Text Rendering Program
Tabs. 5 and 6 show the list of the parameters and their associated feasible sets and variation degrees in the MNIST with Text Rendering Program experiments. The total number of PE iterations is 4. Following [22], we set the number of generated samples to be 60,000.
@@ I.1 MNIST with Text Rendering Program
I.2 CelebA with Generated Images from Computer Graphics-based Render
The variation degrees $\gamma$ across PE iterations are [1000, 500, 200, 100, 50, 20]. The total number of PE iterations is 6. Following [22], we set the number of generated samples to be 60,000.
I.3 CelebA with Rule-based Avatar Generator
@@
These are taken from the input parameters to the library [19]. There is no numerical parameter.
For the experiments with only the simulator, the variation degrees $\beta$ across PE iterations are [0.8, 0.6, 0.4, 0.2, 0.1, 0.08, 0.06]. The total number of PE iterations is 7. Following [22], we set the number of generated samples to be 60,000.
For the experiments with both foundation models and the simulator, we use a total of 5 PE iterations so as to be consistent with the setting in [22]. For the RANDOM_API and the first PE iteration, we use the simulator ($\beta=0.8$). For the next 4 PE iterations, we use the same foundation model as in [46] with variation degrees [96, 94, 92, 90]. Following [22], we set the number of generated samples to be 60,000.
Appendix J Efficiency Evaluation
As simulators could be much cheaper to generate samples than the foundation models, we show in this section that Sim-PE is much more efficient than PE in our experiments.
Note that the only difference between Sim-PE and PE is the RANDOM_API and VARIATION_API. Therefore, we focus on comparing the computation time and peak CPU/GPU memory of the APIs. Tabs. 7, 8 and 9 show that Sim-PE APIs require far less time than PE. For instance, on MNIST, each PE’s API call takes over 2400 GPU seconds, whereas Sim-PE takes less than 30 CPU seconds–an 80x speedup, not to mention the lower cost of CPU than GPU. Consequently, each Sim-PE iteration is much faster than PE (Tab. 10). The only Sim-PE operation requiring GPU is computing the embedding and nearest neighbors of simulator-generated data (§ 3.3), which is a one-time cost per dataset. Even this one-time process is significantly faster than one PE API call. Sim-PE with data access requires more CPU memory to store simulator-generated data, but this can be easily optimized by loading only the needed data.
Time Peak CPU Memory Peak GPU Memory
PE RANDOM_API 3920.30 seconds (GPU) 15292.98 MB 13859.10 MB
VARIATION_API 2422.44 seconds (GPU) 15880.25 MB 16203.89 MB
Sim-PE RANDOM_API 27.17 seconds (CPU) 758.90 MB 0 MB
VARIATION_API 18.51 seconds (CPU) 1083.11 MB 0 MB
Table 7: Efficiency comparison of PE and Sim-PE on the MNIST dataset. Tested on a Linux server with AMD EPYC 7V12 64-Core Processor and one NVIDIA RTX A6000 GPU.
Time Peak CPU Memory Peak GPU Memory
PE RANDOM_API 39272.33 seconds (GPU) 15293.29 MB 13859.10 MB
VARIATION_API 31028.24 seconds (GPU) 15879.47 MB 16203.89 MB
Sim-PE RANDOM_API 0.04 seconds (CPU) 61416.62 MB 0 MB
VARIATION_API 0.03 seconds (CPU) 61427.29 MB 0 MB
Sim-PE Setup 1450.94 seconds (GPU) 61420.69 MB 3728.16 MB
Table 8: Efficiency comparison of PE and Sim-PE (with data access) on the CelebA dataset. Tested on a Linux server with AMD EPYC 7V12 64-Core Processor and one NVIDIA RTX A6000 GPU.
Time Peak Memory Peak GPU Memory
PE RANDOM_API 39272.33 seconds (GPU) 15293.29 MB 13859.10 MB
VARIATION_API 31028.24 seconds (GPU) 15879.47 MB 16203.89 MB
Sim-PE RANDOM_API 46.69 seconds (CPU) 811.72 MB 0 MB
VARIATION_API 48.78 seconds (CPU) 1067.93 MB 0 MB
Table 9: Efficiency comparison of PE and Sim-PE (with code access) on the CelebA dataset. Tested on a Linux server with AMD EPYC 7V12 64-Core Processor and one NVIDIA RTX A6000 GPU.
MNIST CelebA (Sim-PE with data access) CelebA (Sim-PE with code access)
PE 22243.32 279695.52 279695.52
Sim-PE 607.95 441.63 880.38
Table 10: The runtime (seconds) of one iteration in PE and Sim-PE. Tested on a Linux server with AMD EPYC 7V12 64-Core Processor and one NVIDIA RTX A6000 GPU.